Produce the following table:

\frac{e^{x}}{y} \frac{ex}{y}

1 4 5 2 3 4

Statistic

Distribution  Expected value Variance
Binomial(n, p)  np  np(1 − p)
Uniform(α, β)  (β + α)/2  (β − α)^{2}/12
Exponential(λ)  1/λ  1/λ^{2}

Table 1: Means and variances